OpenStudy (anonymous):

If cos t = − 12/13and if the terminal point determined by t is in Quadrant III, find tan t cot t + csc t

OpenStudy (anonymous):

sin t is negative in 3rd quadrant. \[\sin t=\sqrt{1-\cos ^2t}\]\[\cot t=\frac{ 1 }{ \tan t },\tan t \cot t=1,\csc t=\frac{ 1 }{ \sin t }\]